Given the following financial data for company XYZ, calculate their accounts receivable turnover assuming 75% of their sales are on account.

Enter your answer as a number with two places of precision (i.e. 1.23).

xid-100948878_1
Answer #1

accounts receivable turnover = Net credit sales/ Average receivables

Net credit sales = 42000*75% = 31500

Average receivables = (6000 +1500)/2 = 3750  

accounts receivable turnover = 31500/3750 = 8.40 times